# Parameterized sequential functions for temporal properties

### v2

@@ Abstract
Developers of “reactive” [4] software, like operating systems, still need to complete significant code development to have much idea about what performance and correctness problems or overlooked issues will force changes in the design or compromises in design goals. Temporal logic seemed to provide a way for developers to think about and validate these kinds of systems at a higher level, but progress has been slow at best. This paper shows that the kinds of properties that temporal logics promised to address, are among those that can be expressed in ordinary algebra with less notation and without the overhead of formal logic or axiomatic methods, using sequential functions[14] to define ordinary state machines.
@@ 1 Introduction
Temporal logic[11, 10, 7, 3, 5, 6] is based on an intuitively appealing claim that being able to specify when a property must be true can be useful in specifications of computer systems. Examples include properties like: “Always no more than one process can be accessing a mutually exclusive resource”, and “Eventually every scheduled process must start to run”, and “a server cannot send an acknowledgment Until it has received the request.” This paper shows how to express similar and more detailed properties without formal logics or other axiomatic system, using sequential functions [14]. A sequential function is a map from finite sequences of events to outputs so that $f(w)$ is the output of a discrete state system in the state reached by following event sequence $w$ from the initial system state. As seen below, these functions can specify detailed operation of large scale state systems as well as abstract properties and the behavior and architecture of composite and concurrent systems.
Section 2 starts with a short introduction to sequential functions and a definition of temporal-logic-style operators on boolean valued sequential functions so that $(\mathop{\mathit{Always}}P)(w)$, $(\mathop{\mathit{Eventually}}P)(w)$ and so on are well-defined and can be seen to be at least similar to the corresponding temporal qualifiers (which differ among themselves in any case). Section 3 tackles a number of examples, including real-time and concurrent systems, and contrasts specifications that do and do not use the temporal-logic-like operators. Section 4 covers a real-time example used in a paper by Abadi and Lamport[1] and includes a detailed proof. The final section tries for some context.
The goal of this project is perhaps anachronistic: it is to provide human developers of operating systems and other complex “reactive” software [4] with a mathematical basis for specifying and thinking about designs. The reader can form their own opinion about how successful this effort has been and whether it is more or less successful than temporal logic.
@@ 2 Basic methods and Temporal-like operators
Each sequential function is associated with an alphabet – the set of discrete events that can can produce a state change. Every state change is due to some event in the alphabet, although it is common practice to not completely list the alphabet as shown below. State systems are deterministic, so any finite sequence of events determines the state reached by following the sequence from the initial state. All event sequences in this paper are finite.
If $w$ is a event sequence and $f$ is a sequential function, then $f(w)$ is a output in the state reached by following $w$ from the initial state. Following the zero length sequence $\epsilon$ from the initial state leaves the system state unchanged so $f(\epsilon)$ is the output in the initial state. Appending an event $a$ to a sequence, $w\cdot a$, drives the system to the next state so $f(w\cdot a)$ is the output in the state reached after event $a$ drives the system from the state determined by $w$. By convention $\epsilon$ is not an element of any alphabet and $w\cdot\epsilon=w$ leaving the system in the same state (see section 3.3 for why this is convenient).
@@
$f(\epsilon)=c\text{ for constant }c,\ f(w\cdot a)=g(a,f(w)).$
@@
modifies the output and compositions of multiple sequential functions e.g.
@@
define direct products of state systems (which do not interact). More complex products can be defined using sequential functions as transducers: see section 3.3.
Boolean functions have values in the set $\{0,1\}$ with $1$ for true and 0 for false. Boolean expressions like $f(w)<c$ will be treated as abbreviations for characteristic boolean functions like
@@
Sequential functions depend on a single event sequence argument, which is usually given as the first argument. Multiple arguments can mean that the output set $X$ consists of maps (generally not sequential functions), so $f(w,x,y,z)=f(w)(x,y,z)$ or it may be that $f(w)=(f_{1}(w),f_{2}(w),\cdots f_{n}(w))$ and $f(w,i)=(f(w))_{i}=f_{i}(w)$.
More details on sequential functions can be found in appendix B and in the earlier paper [14] where it is shown that sequential functions are essentially just Moore-type state machines and products of these machines. The focus here is on intuition and applications not the underlying automata theory.
@@ 2.1 Temporal style operators
Concatenation of finite sequences, sequence prefixes, and sequence length are all well known operations, but definitions can be found in appendix A in case of doubt.
@@ 2.1.1 Some Conventions
Nested temporal style operators sometimes require a little care – like nested summations or integrals do. Consider $(\mathop{\mathit{Always}}(\mathop{\mathit{Eventually}}P))(w)$. Let $Q(w)=(\mathop{\mathit{Eventually}}P)(w)$. Then the original expression can be written $(\mathop{\mathit{Always}}Q)(w)$ which means that for all $z$ so that $Q(w\mathop{\mathtt{concat}}z)$ is defined, $Q(w\mathop{\mathtt{concat}}z)$. This last expression is defined to mean
@@
Suppose $Sent(w,i,m)$ is true iff network element $i$ has sent message $m$, $Received(w,i,m)$ has the obvious meaning and $Sending(w,i,m)$ is true if $i$ is currently sending the message. For these, suppose that if $Sent(w,i,m)>Sent(w\cdot a,i,m)$ is impossible so that a message never becomes unsent, and the same for $Received$. Then
@@
$\begin{array}[]{l}\text{If }Sent(w,i,m)<Received(w,i,m)\\ \text{ then }(\mathop{\mathit{Eventually}}Sending)(w,i,m))\end{array}$ (2)
@@
Then for some $n\geq 0$ (SinceSending)(w,i,m))(1Sent(w,i,m)))<n(Since\ Sending)(w,i,m))*(1-Sent(w,i,m)))<n is equivalent to 1 above and for some $n_{1}\geq 0$ $((SinceSent)(w,i,m))*(1-Received(w,i,m^{\prime}))<n_{1}$ for some fixed $n_{1}$, is the same constraint as in 2 above. These constraints assume that there are no errors, which is not always a good assumption. It may be better to allow the network element to conclude that a client has failed or is unreachable if for some $k$ (SinceSent)(w,i,m)>k(Since\ Sent)(w,i,m)>k and $Received(w,i,m)=0$. Section 3.4 shows how to count elapsed time instead of the number of events.
@@ 3.0.1 Simple examples
Some simple examples illustrate and will be useful in section 4. The common theme is to define state recursively. The functions “tailtail” and “headhead used below are defined in appendix A. used below.
@@ 3.3 Multiple processes
Suppose the schedule state systems from section 3.1 are components of an operating system or network system with its own event alphabet and there is set $ProcessIds$ of process identifiers. For $p\in ProcessIds$ a set of process identifiers, let:
@@
Connector function definitions have two standard forms. The simpler case, where components advance by one or zero events on every event of the enclosing system looks like this:
$u_{p}(\epsilon)=\epsilon,\text{ and }u_{p}(w\cdot a)=u_{p}(w)\cdot g_{p}(w,a).$
The equation $u_{p}(\epsilon)=\epsilon$ puts the components in their initial states when the operating system is in its initial state. The initial value can be any constant sequence over the component alphabet. The map $g_{p}$ is called the tail function and $g_{p}(w,a)\in A_{sched}\cup\{\epsilon\}$ When the system advances by event $a$ from the state determined by $w$, schedule component $p$ advances by $g_{p}(w,a)$ or remains in its current state if $(g_{p}(w,a)=\epsilon)$. The more general form components can also advance by multiple steps on a single event for the enclosing state system (see appendix B).
There can be many different kinds of components in an interconnected system and figure 3 is a partial illustration of the operating system with components. For the $Req$ sequential function of section 3.1 there might be connector maps $v_{p}$ for each $p\in Processes$ and
@@
This property could be made more concrete with:
$\begin{array}[]{l}\text{If }g_{p}(w,a)=(run,c)\text{ then }\Sigma_{p}(Pstatus(w,p)=(running,c))=0\\ \text{ and }\Sigma_{p}(g_{p}(w,a)=(run,c))=1\\ \text{ and }Pstatus(w,p)=ready\end{array}$
@@ 3.4 Real-time process constraints
$Delay(w\cdot a,p)=\begin{cases}Delay(w,p)+Elapsed(w,a)&\text{if }Pstatus(w\cdot a,p)=ready\\ &\text{and }Pstatus(w,p)=ready\\ 0&\text{if }Pstatus(w\cdot a,p)\neq ready\\ Delay(w,p)&\text{otherwise}\end{cases}$
@@
$\text{for some }x_{0}>0,RTSched(w,x_{0})$
@@ 4 A Message queue
From the outside, ignoring the components for the moment, suppose there is an alphabet $A_{queue}$ which contains events including deqdeq and $(input,m)$ for every message $m\in Msgs$ and other events which can be left for later. The output is a pair of the message at the head of the queue and a bit that alternates on each message.
@@
Here, the buffered queue is constructed from the store, toggle, and queue components from specification 1 above. There are three toggles and two buffers and one transparent queue. A $(input,m)$ event stores message $m$ in the input buffer and toggles the input toggle. A deqdeq event causes the internal queue to move the head of the queue to the output buffer and toggles the output toggle. An internal enqenq event copies the message in the input buffer to the internal queue and flips the $Last$ toggle bit. The components of the queue and connector functions are given below, with component names starting with a capital letter (e.g. $Last)$ and connector functions all lower case: lastlast.
@@
There is a set of messages $\mathit{Msgs}$. InvalInval and $Outval$ are stores over $\mathit{Msgs}$ with initial value nullmnullm; $Q$ is a transparent queue over $\mathit{Msgs}$; $Inb,Last$, and $Outb$ are toggles with initial value 0.
The connector maps are $ov,ob,inb,last,\text{ and }q: invinv$ for $Inval$, inbinb for $Inb$, ovov for $Outval$, obob for $Outb$, lastlast for $Last$, $q$ for $Q$.
The initial values of the connector functions are all $\epsilon$.
$inv(\epsilon)=ov(\epsilon)=inb(\epsilon)=last(\epsilon)=ob(\epsilon)=q(\epsilon)=\epsilon$
The toggles all start out with output 0 per the specification. The message buffers are both set to $\mathit{nullm}$ initially. For example, since $inv(\epsilon)=\epsilon$ and $Invar$ is a store with initial value nullmnullm, $Invar(inv(\epsilon))=nullm$.
Each component advances only one or zero steps for every composite system event in $A_{queue}$ since we let $w\cdot\epsilon=w$.
All connector maps have the form
$u(\epsilon)=c,\ u(w\cdot a)=u(w)\cdot{u}^{\mathsf{T}}(w,a)$
so specifying the recursive tail is most of what is needed. The convention that the tail of $u$ is named ${u}^{\mathsf{T}}$ is just a naming convention, the T doesn’t have any other meaning.
An $(input,m)$ event produces an event stores $m$ to the $Inval$ component and toggles the $Inb$ toggle. Use the convention that $(E\neq E^{\prime})$ has value one if the two expressions are not equal and 0 otherwise and then set
$WaitEnq(w)=(Line(line(w))\neq Inb(inb(w)))$
The intent of the protocol with the toggles is that $WaitEnq(w)=1$ when $Inval(iv(w))$ is an input message that has not yet been queued on the internal queue. The most complex connection map is $q$ which controls events for the input queue and which is not completely specified but is constrained from enqueuing unless the $Last$ and $Inb$ values differ. Incomplete specification takes the place of nondeterminism.
@@
Finally:
$Queue(w)=(Oval(ov(w)),Obit(ob(w)))$
@@ 4.1 Real-time
To patch up the poor design of the protocol in this queue, Abadi and Lamport propose to put timing constraints on input and enq operations. The basic idea here is that when an inputinput event happens, a timer starts and the matching enqenq must happen before $t_{0}$ seconds pass, but the next input event cannot happen until more than $t_{0}$ seconds have passed. It’s pretty obvious that these timing constraints fix the problem with dropped messages, but see section 4.2.
@@
There cannot be an inputinput event without triggering a timing error until $SinceI(w)>t_{0}$. There must be an enqenq event before $SinceI(w)\geq t_{0}$.
@@
The timing specification is not requiring that events follow the specification which could be done with $\mathop{\mathit{Always}}Tspec(w)=1$. Instead the specification allows us to only consider states that are reached without violations of the specification.
For comparison, see page 1553 of [1] which is similar at least in intent.
@@ 4.2 Correctness
$Q_{O}(w\cdot a)=\begin{cases}Q_{O}(w)\cdot m&\text{if }{q}^{\mathsf{T}}(w,a)=(deq,m)\text{ for some }m\in\mathit{Msgs}\\ Q_{O}(w)&\text{otherwise}\end{cases}$
@@
$\text{If }WaitEnq(w)=0\text{ then }Q_{I}(w)=Q_{O}(w)\mathop{\mathtt{concat}}Q(q(w))$
$\text{If }WaitEnq(w)=1\text{ then }Q_{I}(w)=Q_{O}(w)\mathop{\mathtt{concat}}Q(q(w))\cdot Inval(in(w))$
The difference is that when $WaitEnq(w)=1$ the value in the input buffer is the last item that has shown up on the input queue, otherwise the input buffer can be ignored. It’s an immediate consequence of this theorem that $Q_{O}(w)$ is a prefix of $Q_{I}(w)$.
The proof of 1 is by induction on the event sequence. Initially, when $w=\epsilon$, all the queues are empty and $WaitEnq(\epsilon)=0$ so there is nothing to prove. Suppose the theorem is true for $w$ and consider $w\cdot a$. We assume $Tspec(w\cdot a)$ otherwise there is nothing to prove. There are four cases.
@@
    By the definition of $WaitEnq$, one of the two toggles has to change thanks to $a$. If ${q}^{\mathsf{T}}(w,a)=(deq,m)$ that would change the $Last$ toggle but because $WaitEnq(w)=0$ ${q}^{\mathsf{T}}(w,a)\neq(deq,m)$ for any $m$ so it must be that $a=(input,m)$ for some $m$ and $Inb$ changes. By the inductive assumption and because $WaitEnq(w)=0$ $Q_{I}(w)=Q_{O}(w)\mathop{\mathtt{concat}}Q(q(w))$. Because $a=(input,m)$, $Q_{I}(w\cdot a)=Q_{I}(w)\cdot m$. By the definition of ${inv}^{\mathsf{T}}$ and because $a=(input,m)$, $Inval(inv(w\cdot a))=m$. We know $a\neq deq$ so ${q}^{\mathsf{T}}(w,a)\neq deq$ – so $Q(q(w))=Q(q(w\cdot a))$. This proves the first case.
@@
    In this case $a\neq(input,m)$ because otherwise $WaitEnq(w\cdot a)=1$ unless the $Last$ toggle flips too. But $Last$ cannot change because that needs ${q}^{\mathsf{T}}(w,a)=(enq,m)$ and that cannot be true if $WaitEnq(w)=0$. So $Q_{I}(w\cdot a)=Q_{I}(w))$. If $a=deq$ then if $Q(q(w))=\epsilon$ nothing happens to $Q_{O}$ or $Q$. If $Q(q(w))\neq\epsilon$ and $a=deq$, then ${q}^{\mathsf{T}}(w,a)=deq$ so $Q(q(w\cdot a)=Tail(Q(w))$ and $Q_{O}(w\cdot a)=Q_{O}(w)\cdot Head(Q(q(w)))$ but
    $Q_{O}(w)\cdot Head(Q(q(w))\mathop{\mathtt{concat}}Tail(Q(q(w))$
@@
    $Q_{O}(w)\mathop{\mathtt{concat}}Q(q(w)).$
    Otherwise ${q}^{\mathsf{T}}(w,a)=\epsilon$ That takes care of this case.
@@
    Because $WaitEnq$ remains true, ${q}^{\mathsf{T}}(w,a)\neq(enq,m)$ for any $m$ so no new elements are added to $Q(q(w\cdot a))$. Because we assume $Tspec(w\cdot a)$, $WaitEnq(w)*SinceI(w)<t_{0}$ so $a$ is not an input event so ${inv}^{\mathsf{T}}(w,a)=\epsilon$ and $Inval(inv(w\cdot a))=Inval(inv(w))$. If $a\neq deq$, it must be that ${q}^{\mathsf{T}}(w,a)=\epsilon$ so $Q(q(w\cdot a))=Q(q(w))$ when $a\neq deq$ and similarly $Q_{O}(w\cdot a)=Q_{O}(w)$ because ${ov}^{\mathsf{T}}(w,a)=\epsilon$. If $a=deq$ then than $Q(q(w\cdot a))=Tail(Q(q(w)))$ and $Q_{O}(w\cdot a)=Q_{O}(w)\cdot Head(Q(q(w)))$ so the constructed queue is unchanged.
@@
    $Q_{I}(w)=Q_{O}(w)\mathop{\mathtt{concat}}Q(q(w))\cdot Inv(inv(w).$
    Either $SinceI(w)<t_{0}$ or $TSpec(w)=0$. We’re assuming the second is false, so by the same reasoning $a\neq(input,m)$ (or else $Tspec(w\cdot a)=0$). That means that $q(w,a)=(enq,Inv(inv(w)))$ to make $WaitEnq(w\cdot a)=0$. As a result
    $Q(q(w\cdot a))=Q(q(w))\cdot Inv(inv(w)).$

### v3

@@ Abstract
Temporal logic is based on an intuitively appealing claim that being able to specify when a property will be true can be useful in specifications of computer software and hardware - particularly in reactive systems [6]. This paper shows that temporal properties are among the properties that can be expressed directly and more simply in terms of state machines using sequential functions[17] to define state machine behavior and compositional architecture. Specifications and proofs can be carried out in ordinary algebra without the need for a formal language or axiomatic proof system. The compositional properties of sequential functions also simplify specification of composite and concurrent systems.
@@ 1 Introduction
Temporal logic[13, 12, 10, 3, 7, 9] is based on an intuitively appealing claim that being able to specify when a property must be true can be useful in specifications of computer systems. Examples include properties like: “Always no more than one process can be accessing a mutually exclusive resource”, and “Eventually every scheduled process must start to run”, and “a server cannot send an acknowledgment Until it has received the request.” This paper shows how to express similar and more detailed properties without formal logics or other axiomatic system using sequential functions [17]. A sequential function is a map from finite sequences of events to outputs so that $f(w)$ is the output of a discrete state system in the state reached by following event sequence $w$ from the initial state. As seen below, these functions can specify detailed operation of large scale state systems as well as abstract properties and the behavior and architecture of composite and concurrent systems.
Section 2 starts with a short introduction to sequential functions and a definition of temporal-logic-style operators on boolean valued sequential functions so that $(\mathop{\mathit{Always}}P)(w)$, $(\mathop{\mathit{Eventually}}P)(w)$ and so on are well-defined and can be seen to be at least similar to the corresponding temporal qualifiers (which differ among themselves in any case). Section 3 tackles a number of examples, including real-time and concurrent systems, and contrasts specifications that do and do not use the temporal-logic-like operators. Section 4 covers a real-time example used in a paper by Abadi and Lamport[1] but unlike that paper includes a complete proof. The final section tries for some context.
The goal of this project is perhaps anachronistic: it is to provide human developers of operating systems and other complex “reactive” software [6] with a mathematical basis for specifying and thinking about designs. The reader can form their own opinion about how successful this effort has been.
@@ 2 Basic methods and Temporal-like operators
Each sequential function is associated with an alphabet – the set of discrete symbols representing the events that can can produce a state change. State systems are deterministic, so any finite sequence of events determines the state reached by following the sequence from the initial state. Event sequences in this paper are always finite. Much of the expressive power of sequential functions derives from how convenient it is to make partial specifications: instead of making nondeterminism an intrinsic property, we can just leave some behavior unspecified.
Following the zero length sequence $\epsilon$ from the initial state leaves the system state unchanged so for sequential function $f$, $f(\epsilon)$ is the output in the initial state. Appending an event $a$ to a sequence, $w\cdot a$, drives the system to the next state so $f(w\cdot a)$ is the output in the state reached after event $a$ drives the system one event from the state determined by $w$.
@@
 $\text{ for constant }c,\ f(\epsilon)=c,\text{ and }f(w\cdot a)=g(a,f(w)).$
@@
modifies the output. Compositions of multiple sequential functions e.g.
@@
define direct products of state systems (which do not interact). An analog of simultaneous recursion of arithmetic primitive recursive functions [11] defines a system of interacting components:
$f(w)=(f_{1}(w_{1}),\dots f_{n}(w_{n}))$
where each $w_{i}=w_{i}(w)$ is a primitive recursive sequential function that determines component interactions. See section 3.3 for details.
Boolean valued sequential functions that have values in the set $\{0,1\}$ with $1$ for true and 0 for false can be used to specify properties. Boolean expressions like $f(w)<c$ will be treated as abbreviations for characteristic boolean functions like
@@
Sequential functions depend on a single event sequence argument, which is usually given as the first argument. Multiple arguments can mean either that that the output set $X$ consists of maps (generally not sequential functions), so $f(w,x,y,z)=(f(w))(x,y,z)$.
The focus here is on intuition and applications not the underlying automata theory.
@@ 2.1 Temporal style operators
Sequence prefixes and sequence length are well known operations, but definitions can be found in appendix A in case of doubt.
@@ 2.1.1 Some Conventions
Nested temporal style operators sometimes require a little care – like nested summations or integrals do. Consider $(\mathop{\mathit{Always}}(\mathop{\mathit{Eventually}}P))(w)$. Let $Q(w)=(\mathop{\mathit{Eventually}}P)(w)$. Then the original expression can be written $(\mathop{\mathit{Always}}Q)(w)$ which means that for all $z$ so that $Q(w\mathop{\mathtt{concat}}z)$ is defined, $Q(w\mathop{\mathtt{concat}}z)$. Expanding $Q$ the last expression means
@@
Suppose $Sent(w,i,m)$ is true iff network element $i$ has sent message $m$, $Received(w,i,m)$ has the obvious meaning and $Sending(w,i,m)$ is true if $i$ is currently sending the message. For these, suppose that $Sent(w,i,m)>Sent(w\cdot a,i,m)$ is impossible so that a message never becomes unsent, and the same for $Received$. Then
@@
$\begin{array}[]{l}\text{If }Sent(w,i,m)>Received(w,i,m^{\prime})\\ \text{ then }(\mathop{\mathit{Eventually}}Sending)(w,i,m))\end{array}$ (2)
@@
Then for some $n\geq 0$, (SinceSending)(w,i,m))(1Sent(w,i,m)))<n(Since\ Sending)(w,i,m))*(1-Sent(w,i,m)))<n is equivalent to 1 above and for some $n_{1}\geq 0$ $((SinceSent)(w,i,m))*(1-Received(w,i,m^{\prime}))<n_{1}$ for some fixed $n_{1}$, is the same constraint as in 2 above. These constraints assume that there are no errors, which is not always a good assumption. It may be better to allow the network element to conclude that a client has failed or is unreachable if for some $k$ (SinceSent)(w,i,m)>k(Since\ Sent)(w,i,m)>k and $Received(w,i,m)=0$. Section 3.4 shows how to count elapsed time instead of the number of events.
@@ 3.0.1 Simple examples
Some simple examples illustrate and will be useful in section 4. The common theme is to define state recursively. The functions “tailtail” and “headhead used below are as usual and are defined in appendix A in case of doubt.
@@ 3.3 Multiple processes
Suppose the schedule state systems from section 3.1 are components of an operating system or network system with its own event alphabet and there is set $ProcessIds$ of process identifiers. For $p\in ProcessIds$, let:
@@
Connector functions are defined recursively:
$u_{p}(\epsilon)=\epsilon,\text{ and }u_{p}(w\cdot a)=u_{p}(w)\mathop{\mathtt{concat}}G_{p}(w,a).$
The equation $u_{p}(\epsilon)=\epsilon$ puts the process components in their initial states when the operating system is in its initial state. The initial value can be any constant sequence over the component alphabet. The map $G_{p}$ is called the tail function and $G_{p}(w,a)$ is what happens to the component $p$ when the system advances by an $a$ event from the system state determined by $w$. If $G_{p}(w,a)=\epsilon$, component $p$ remains in the same state. For this simple example suppose each process component only advances 0 or $1$ steps on each step of the enclosing system and write $a$ for $\langle a\rangle$ if it is clear from the context that the single element sequence being used.
There can be many different kinds of components in an interconnected system and figure 3 is a partial illustration of the operating system with components. For the $Req$ sequential function of section 3.1 there might be connector maps $V_{p}$ for each $p\in Processes$.
@@
This property could be made more direct with:
$\begin{array}[]{l}\text{If }G_{p}(w,a)={(run,c)}\text{ then }Pstatus(w,p)=ready\\ \text{ and }\Sigma_{p}(Pstatus(w,p)=(running,c))=0\\ \text{ and }\Sigma_{p}(G_{p}(w,a)={(run,c))}=1\end{array}$
@@ 3.4 Real-time process constraints
$Delay(w\cdot a,p)=\begin{cases}Delay(w,p)+Elapsed(w,a)&\text{if }Pstatus(w\cdot a,p)=ready\\ &\text{and }Pstatus(w,p)=ready\\ 0&\text{otherwise}\end{cases}$
@@
$\text{for some }x,\ RTSched(w,x)$
@@ 4 A Message queue
From the outside, ignoring the components for the moment, suppose there is an alphabet $A_{queue}$ which contains events including deqdeq and $(input,m)$ for every message $m\in\mathit{Msgs}$ and other events which can be left unspecified. The output is a pair of the message at the head of the queue and a bit that alternates on each message.
@@
The queue is constructed from the store, toggle, and queue components from specification 1 above.
@@
$\mathit{Msgs}$ is a set of messages.
InvalInval and $Outval$ are stores over $\mathit{Msgs}$ with initial value nullmnullm;
InbInb, OutbOutb, and $Last$ are toggles with the same initial value.
$Q$ is a transparent queue over $\mathit{Msgs}$;
A $(input,m)$ event stores message $m$ in the input buffer and toggles the input toggle. A deqdeq event causes the internal queue to move the head of the queue to the output buffer and toggles the output toggle. An internal enqenq event copies the message in the input buffer to the internal queue and flips the $Last$ toggle bit. The intent of the protocol with the toggles is that the $Inb$ and $Last$ toggles are not equal if and only if there is a live message in $Inval$ tht has not yet been enqued in the internal queue.
The connector maps are $w_{1}$ for $Inb$ and $Inval$, $w_{2}$ for the toggle $Last$, $w_{3}$ for the internal queue $Q$, and $w_{4}$ for the output bit toggle and output buffer. Say $w_{i}(w\cdot a)$ adds an event $a$ if and only if $w_{i}(w\cdot a)=w_{i}(w)\cdot a$. Say $w_{i}(w\cdot a)$ adds nothing if and only if $w_{i}(w\cdot a)=w_{i}(w)$. The specification for $w_{3}$ is different from the others because we don’t want to fully specify when $w_{3}(w\cdot a)$ adds an enqenq event. In the Abadi and Lamport specification this is a nondeterministic choice, here it depends on the unspecified variations among the components, the other input events, and possibly other factors not specified. Section 4.1 adds real-time constraints.
@@
The queue device itself then just selects from the tuple value of $RawDevice$.
Specification 4
$Queue(w)=((RawDevice(w))_{5},(RawDevice(w))_{6})$
@@ 4.1 Real-time
To patch up the poor design of the protocol in this queue, Abadi and Lamport propose to put timing constraints on input and enq operations. The basic idea here is that when an $(input,m)$ event happens, a timer starts and the matching enqenq must happen before $t_{0}$ seconds pass, but the next input event cannot happen until more than $t_{0}$ seconds have passed. It’s pretty obvious that these timing constraints fix the problem with dropped messages, but see section 4.2.
@@
There cannot be an $(input,m)$ event without triggering a timing error until $SinceI(w)>t_{0}$. But also once there is an input event, there must be an enqenq event before $SinceI(w)\geq t_{0}$.
To make the specification more compact: $WaitEnq$ indicates if there is a live message in the inbuffer according to the protocol.
$WaitEnq(w)=\begin{cases}1&\mbox{ if }Last(w_{2}(w)))\neq Inb(w_{1}(w)))\\ 0&\text{otherwise}\end{cases}$
@@
The timing specification is not requiring that events follow the specification which could be done with $\mathop{\mathit{Always}}Tspec(w)=1$. Instead the specification allows us to distinguish states that are reached without violations of the specification. Those reached otherwise have unspecified results.
For comparison, see page 1553 of [1] which is similar in intent.
@@ 4.2 Correctness
$Q_{O}(w\cdot a)=\begin{cases}Q_{O}(w)\cdot m&\text{if }w_{3}(w\cdot a)=w_{3}(w)\cdot(deq,m)\text{ for some }m\in\mathit{Msgs}\\ Q_{O}(w)&\text{otherwise}\end{cases}$
@@
$\text{If }WaitEnq(w)=0\text{ then }Q_{I}(w)=Q_{O}(w)\mathop{\mathtt{concat}}Q(w_{3}(w))$
$\text{If }WaitEnq(w)=1\text{ then }Q_{I}(w)=Q_{O}(w)\mathop{\mathtt{concat}}Q(w_{3}(w))\cdot Inval(in(w))$
In either case, $Q_{O}(w)$ is a prefix of $Q_{I}(w)$. The difference is that when $WaitEnq(w)=1$ the value in the input buffer is the last item that has shown up on the input queue, otherwise the input buffer can be ignored. This theorem also shows that the protocol works as long at $Tspec(w)$ remains true.
The proof of 1 is by induction on the event sequence. The proof here is in too much detail, just to show it can be done completely. Initially, when $w=\epsilon$, all the queues are empty and $WaitEnq(\epsilon)=0$ so there is nothing to prove. Suppose the theorem is true for $w$ and consider $w\cdot a$. We assume $Tspec(w\cdot a)$ otherwise there is nothing to prove. There are four cases depending on the four possible values of $(WaitEnq(w),WaitEng(w\cdot a))$.
@@
    By the definition of $WaitEnq$, one of the two toggles has to change thanks to $a$. Because $WaitEnq(w)=0)$, $w_{3}(w\cdot a)$ cannot add a $(deq,m)$ event that would change the $Last$ toggle. So the alternative must be that $a=(input,m)$ for some $m$. $m$ and $Inb$ changes. By the inductive assumption and because $WaitEnq(w)=0$ $Q_{I}(w)=Q_{O}(w)\mathop{\mathtt{concat}}Q(w_{3})$. Because $a=(input,m)$, $Q_{I}(w\cdot a)=Q_{I}(w)\cdot m$. By the definition of $w_{1}$, it must be that $Inval(w\cdot a)=m$ We know $a\neq deq$ so $w_{3}(w\cdot a)$ cannot add deqdeq – so $Q(w_{3}(w))=Q(w_{3}(w\cdot a))$. It follows that $Q_{I}(w\cdot a)=Q_{O}(w\cdot a)\mathop{\mathtt{concat}}Q(w_{3}(w\cdot a)\cdot Inval(w_{1}(w\cdot a)$.
@@
    In this case $a\neq(input,m)$ because otherwise $WaitEnq(w\cdot a)=1$ unless the $Last$ toggle flips too. But $Last$ cannot change because $WaitEnq(w)=0$ means that $w_{3}(w\cdot a)$ does not add an $(enq,m)$ which is required to toggle $Last$. Because $a\neq(input,m)$ it follows that $Q_{I}(w\cdot a))=Q_{I}(w))$. If $a=deq$ then if $Q(w_{3}(w))=\epsilon$ nothing happens to $Q_{I}$, $Q_{O}$ or $Q$. If $Q(w_{3}(w))\neq\epsilon$ and $a=deq$, then $w_{3}(w\cdot a)=w_{3}(w)\cdot deq$ so $Q(w_{3}(w\cdot a)=Tail(Q(w_{3}(w)))$ and $Q_{O}(w\cdot a)=Q_{O}(w)\cdot Head(Q(w_{3}(w)))$ but
    $Q_{O}(w)\cdot Head(Q(w_{3}(w))\mathop{\mathtt{concat}}Tail(Q(w_{3}(w))$
@@
    $Q_{O}(w)\mathop{\mathtt{concat}}Q(w_{3}(w)).$
    Otherwise $w_{3}(w\cdot a)$ adds $\epsilon$ and nothing changes. That takes care of this case.
@@
    Because $WaitEnq$ remains true, $w_{3}(w\cdot a)$ does not add $(enq,m)$ for any $m$ so no new elements are added to $Q(w_{3}(w\cdot a))$. Because we assume $Tspec(w\cdot a)$, $WaitEnq(w)*SinceI(w)<t_{0}$ so $a$ is not an $(input,m)$ for any $m$ and thus $w_{1}(w\cdot a)=w_{1}(w)$ and $Inval(inv(w\cdot a))=Inval(inv(w))$. If $a\neq deq$, it must be that $w_{3}(w\cdot a)$ adds $\epsilon$ so $Q(w_{3}(w\cdot a))=Q(w_{3}(w))$ and similarly $Q_{O}(w\cdot a)=Q_{O}(w)$ because $w_{4}(w\cdot a)i=w_{4}(w)$. If $a=deq$ then then $Q(w_{3}(w\cdot a))=Tail(Q(w_{3}(w)))$ and $Q_{O}(w\cdot a)=Q_{O}(w)\cdot Head(Q(w_{3}(w)))$ so the constructed queue is unchanged.
@@
    $Q_{I}(w)=Q_{O}(w)\mathop{\mathtt{concat}}Q(w_{3}(w))\cdot Inv(inv(w).$
    Either $SinceI(w)<t_{0}$ or $TSpec(w)=0$. We’re assuming the second is false, so by the same reasoning $a\neq(input,m)$ (or else $Tspec(w\cdot a)=0$). That means that $w_{3}(w\cdot a)$ adds $(enq,Inv(w_{1}(w)))$ to make $WaitEnq(w\cdot a)=0$. As a result
    $Q(w_{3}(w\cdot a))=Q(w_{3}(w))\cdot Inv(w_{1}(w)).$